The equation of a line through the point Q ( 1,2) whose distance from the point P (3,1) is greatest , is 
 

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
P(1,2) , Q(3,1)
Slope of PQ = 1-2 / 3-1 = -1/2 
Slope of line  perpendicular to this is 2 
y-2 = 2x-2 
y = 2x 
Hope this helps